# Phase locking and multistability in the topological Kuramoto model on cell complexes

### v1

@@ Abstract
The topological Kuramoto model generalizes classical synchronization models by including higher-order interactions, with oscillator dynamics defined on cells of arbitrary dimension within simplicial or cell complexes. In this article, we demonstrate multistability in the topological Kuramoto model and develop the topological nonlinear Kirchhoff conditions algorithm to identify all phase-locked states on arbitrary cell complexes. The algorithm is based on a generalization of Kirchhoff’s laws to cell complexes of arbitrary dimension and nonlinear interactions between cells. By applying this framework to rings, Platonic solids, and simplexes, as minimal representative motifs of larger networks, we derive explicit bounds (based on winding number constraints) that determine the number of coexisting stable states. We uncover structural cascades of multistability, inherited from both lower and higher dimensions and demonstrate that cell complexes can generate richer multistability patterns than simplicial complexes of the same dimension. Moreover, we find that multistability patterns in cell complexes appear to be determined by the number of boundary cells, hinting a possible universal pattern.
@@ I Introduction
From neuronal activity to power grid operation, synchronization phenomena are a hallmark of many natural and engineered systems [1, 2]. For such systems, the Kuramoto model of coupled oscillators has served as a prototypical model for understanding how synchronization emerges from local interactions [3, 4]. While early studies focused on globally coupled systems [5], later work extended the model to complex networks, revealing that network topology plays a crucial role in determining the onset and nature of synchronization [6]. Such systems often exhibit multistability, where multiple distinct synchronized states can coexist [7, 8, 9, 10].
@@
In their classical form, the coupling between the oscillators is pairwise. This assumption is valid for many systems, but in various contexts, such as neuronal assemblies [11], the brain connectome [12, 13, 14] multimode lasers, [15], or protein interaction networks [16], it is likely too restrictive to capture all the relevant couplings in the system. Group interactions, encoded by higher-order connectivity structures, are increasingly recognized as being crucial to many collective dynamics [17, 18, 19, 20, 21, 22]. A variety of recent advances in network science [18, 19, 23, 12, 24, 25, 13, 26, 27, 28, 14] have introduced models where the dynamics is not defined only on nodes, but also lives on edges, faces, or higher-dimensional elements of simplicial or cell complexes. The topological Kuramoto model [29, 30, 31, 32] is among the most prominent dynamical models in this framework, revealing universal features absent in conventional networks, such as abrupt synchronization transitions  [33, 34] or topologically-induced synchronization obstructions  [35, 36]. While some aspects of multistability [37, 38] and phase locking [32] in the topological Kuramoto model have been analyzed separately, there is a lack of systematic frameworks to analyse how multistable phase-locked states can arise from higher-order interactions.
@@
To bridge this gap, here we investigate phase locking and multistability in the topological Kuramoto model defined on general cell complexes. Unlike earlier works which employ the master stability function formalism to study synchronization in this model [36, 39, 40], we propose the topological nonlinear Kirchhoff conditions algorithm, which may be seen as a generalization of Kirchhoff’s current and voltage laws. Moreover, thus far most studies have focused on simplicial complexes. Yet many real-world structures, such as architectural trusses, social interactions or protein complexes, are better described as cell complexes [41, 42, 43, 44], which consist of larger polyhedral faces. Despite their prevalence, the study of dynamical phenomena on cell complexes is still in its infancy.
@@
Our novel algorithm systematically identifies distinct phase-synchronized states in cell complexes, and allows us to define rigorous criteria for when going beyond pairwise interactions pays off in terms of multistability. We demonstrate our approach on three minimal but paradigmatic classes of cell complexes: (i) rings, (ii) Platonic solids, and (iii) simplexes. These basic motifs can be combined to build larger complexes with potentially even richer multistability. Our framework can be applied to any cell complex and any dimension. For simplicity, however, we focus here on dynamics supported on edges, with interactions mediated through nodes and faces (2-simplices). This case is moreover particularly relevant to applications (see, e.g., Ref. [43] and references therein).
Our main findings are:
  • Cell complexes can support richer multistability than simplicial complexes of the same dimension. This extends previous results on their enhanced synchronizability [36].
  • Higher-order connectivity can result in structural cascades of multistability, which involve sequences of stable states inherited from both lower and higher dimensions. The number of stable states increases with system size and arises from winding-number bounds determined by the cell boundaries.
  • We find indications of possible universality patterns: different complexes, such as dual Platonic solids, have the same multistability patterns if they share certain structure-dependent parameters, related to the number of boundary cells in the upper and lower dimension.
Our results show how higher-order topology governs synchronization and the emergence of multistability, revealing new mechanisms for collective behavior in complex systems with non-pairwise interactions.
@@ II Cell complexes
Cell complexes are generalised network structures that contain cells of different dimensions [41, 12, 42]. In addition to nodes $(n=0)$ and edges ($n=1$) found in standard graphs, they can include polygons $(n=2)$, polytopes $(n=3)$ and other higher-dimensional elements. Most studies on synchronization focus on the special case of simplicial complexes  [29, 32], where each cell is an Euclidean simplex, but many real-world systems are better described by cell complexes.
@@
Figure 1: Two examples of cell complexes for $n=1$. Nodes are colored in black, edges in blue and faces in light green.
@@
Using this property, one can show that every vector ${\boldsymbol{x}}_{[n]}\in\mathbb{R}^{N_{[n]}}$ can be decomposed according to the Hodge decomposition [45] as
@@ III Phase locking in the topological Kuramoto model
In this article we consider a generalized topological Kuramoto model allowing for inhomogeneous coupling constants. The basic dynamic variables are the phases $\theta_{i}(t)$ associated to each $n$-dimensional cell. These phases interact via cells of lower and higher dimension. In the case $n=1$ for example, phases are defined for each edge, and interact via vertices and faces. Summarizing the dynamical variables in the vector ${\boldsymbol{\theta}}=\left(\theta_{1},\ldots\theta_{N_{[n]}}\right)^{\top}$, the equations of motion read
@@
We can decouple the dynamics by multiplying the equations of motion by ${\boldsymbol{B}}_{[n]}$ and ${\boldsymbol{B}}_{[n+1]}^{\top}$ from the left [29]. Defining the variables
@@
In general, a stationary state is linearly stable if the real part of all eigenvalues is negative [46]. If the stationary state satisfies the conditions
@@
we find that all eigenvalues are real and satisfy the following conditions: (i) If an eigenvector $\boldsymbol{v}$} is purely harmonic, then the associated eigenvalue is given by $\lambda=0$. (ii) If an eigenvector $\boldsymbol{v}$} is not purely harmonic, then the associated eigenvalue satisfies $\lambda<0$. A proof is given in appendix A. We thus conclude that a stationary state that satisfies Eq. (11) is linearly stable with respect to perturbations that are orthogonal to the harmonic subspace and neutrally stable with respect to harmonic perturbations. Since harmonic perturbations do not affect the projected quantities $\boldsymbol{\theta}$$^{[+]}$ and $\boldsymbol{\theta}$$^{[-]}$, phase locking is linearly stable.
In the following we will mainly concentrate on phase locked states that satisfy the conditions in Eq. (11) and refer to them as normal fixed points. Stationary states that violate Eq. (11) are typically, but not always, linearly unstable. For the ordinary Kuramoto model $n=0$, it has been shown that stable phase locked states that violate these conditions exist at the border of the stability region [47]. We will return to this question in Sec.V, Fig.2, in the context of polygonal face complexes, where we find such stable phase-locked states that aren’t normal fixed points.
IV Phase locking and multistability
@@
In the case $n=1$, the original dynamical variables $\boldsymbol{\theta}$} are defined on the edges of the cell complex, while the variables $\boldsymbol{\psi}$$^{[+]}$ are defined on the faces, and the variables $\boldsymbol{\psi}$$^{[-]}$ on the nodes. In the case $n=0$, the original dynamical variables $\boldsymbol{\theta}$} are defined on the nodes of the graph, while the auxiliary variables $\boldsymbol{\psi}$$^{[+]}$ are defined on its edges. In the theory of supply networks, the entries of ${\boldsymbol{K}}_{[1]}{\boldsymbol{\psi}}^{[+]}$ are interpreted as flows [48, 49, 50].
@@
The complexity of finding a stationary state is then shifted to the question of whether a solution candidate (16) exists that can be expressed as (12). To approach this topic, we first discuss how to recover the original variables $\boldsymbol{\theta}$$^{[\pm]}$ from the $\boldsymbol{\psi}$$^{[\pm]}$. If the equations Eq. (12) have solutions, they can have the form $\arcsin(\cdot)$ or $\pi-\arcsin(\cdot)$. To keep track of the two options, we partition the set of cells as follows:
$S_{[n+1]}^{\bullet}$ $=\bigg\{j\in S_{[n+1]}\bigg|\cos\left(\theta^{[+]}_{j}\right)>0\bigg\}$
@@
$S_{[n-1]}^{\bullet}$ $=\bigg\{i\in S_{[n-1]}\bigg|\cos\left(\theta^{[-]}_{i}\right)>0\bigg\}$
[…]
We can now consistently express $\boldsymbol{\theta}$$^{[\pm]}$ in terms of $\boldsymbol{\psi}$$^{[\pm]}$, having in mind the definitions in Eq. (12),
$j\in S_{[n+1]}^{\bullet}:\;$ $\left({\boldsymbol{B}}_{[n+1]}^{\top}{\boldsymbol{\theta}}\right)_{j}=f_{+}\left(\psi_{j}^{[+]}\right)=\arcsin\left(\psi_{j}^{[+]}\right)$
$j\in S_{[n+1]}^{\circ}:\;$ $\left({\boldsymbol{B}}_{[n+1]}^{\top}{\boldsymbol{\theta}}\right)_{j}=f_{-}\left(\psi_{j}^{[+]}\right)=\pi-\arcsin\left(\psi_{j}^{[+]}\right)$
$i\in S_{[n-1]}^{\bullet}:\;$ $\left({\boldsymbol{B}}_{[n]}{\boldsymbol{\theta}}\right)_{i}=f_{+}\left(\psi_{i}^{[-]}\right)=\arcsin\left(\psi_{i}^{[-]}\right)$
$i\in S_{[n-1]}^{\circ}:\;$ $\left({\boldsymbol{B}}_{[n]}{\boldsymbol{\theta}}\right)_{i}=f_{-}\left(\psi_{i}^{[-]}\right)=\pi-\arcsin\left(\psi_{i}^{[-]}\right)$ (20)
up to integer multiples of $2\pi$. We summarize these components in the two vector-valued functions ${\boldsymbol{f}}_{[n+1]}({\boldsymbol{\psi}}^{[+]})$ and ${\boldsymbol{f}}_{[n-1]}({\boldsymbol{\psi}}^{[-]})$.
@@
where $\boldsymbol{z}$$^{[\pm]}$ is a vector of integers. In the special case where the kernel is parametrized by a single basis vector, $z^{[\pm]}$ is a scalar. In the case of ordinary graphs, the vectors $\boldsymbol{z}$$^{[\pm]}$ are referred to as winding vectors and its components as winding numbers [8, 7, 9]. We will adopt this notation in the following and discuss the interpretation below.
@@
We briefly discuss three important aspects of this algorithmic approach to phase locking. First, phase locked states with $S^{\circ}=\emptyset$ are always linearly stable (see Sec. III), and as such are especially relevant for applications. States with $S^{\circ}\neq\emptyset$ are typically, but not always unstable. In fact, we will demonstrate the existence of stable phase locked states with $S^{\circ}\neq\emptyset$ close to bifurcations.
Second, the algorithm can be viewed as a topological, nonlinear generalization of Kirchhoff’s circuit laws. This is most easily see for the case of ordinary graphs $n=0$. Here, the entries of ${\boldsymbol{K}}_{[1]}{\boldsymbol{\psi}}^{[+]}$ are interpreted as flows or currents defined on edges [48, 49, 50]. The set of solution candidates obtained by Eq. (16) correspond to all states that satisfy Kirchhoff’s current law at each node. These states can be decomposed into a special solution plus a set of loop flows. If we linearize the sine and arcsine functions and set ${\boldsymbol{z}}^{[\pm]}={\boldsymbol{0}}$, then conditions (22) correspond to Kirchhoff’s voltage law. Together, Kirchhoff’s voltage and current laws then uniquely determine the currents.
Algorithm 1 introduces two generalizations: (i) The nonlinearity of the sine introduces multistability. Since $\sin(\alpha)=\sin(\alpha+2\pi z)$ for all $z\in\mathbb{Z}$, we can obtain different values for phase variables when inverting the sine function. Hence, we can obtain different solutions parametrized by the winding numbers $z$ [8, 7, 9]. (ii) The algorithm generalizes the nonlinear Kirchhoff conditions to cell complexes of arbitrary dimensions.
Third, the possible values of the entries of the vectors $\boldsymbol{z}$$^{[\pm]}$ and thus the number of steady states is bounded. To derive an upper bound, we rewrite Eq. (22) in components and use that $|f_{\pm}|\leq\pi/2$ and obtain
@@
[…]
We recall that the matrices $\boldsymbol{C}$} are closely related to the boundary matrices. We can choose a basis of the kernel such that the $k$th column of ${\boldsymbol{C}}_{[n+1]}$ (${\boldsymbol{C}}_{[n]}^{\top}$) is equal to the $k$th column of ${\boldsymbol{B}}_{[n+2]}$ (${\boldsymbol{B}}_{[n-1]}^{\top}$) unless $k$ corresponds to the harmonic component. A further simplification is possible by noting that the entries of the boundary matrix have the values 1,0,+1-1,0,+1. We then define $m_{k}^{[+]}$ as the number of non-zero elements of the $k$th column of ${\boldsymbol{B}}_{[n+2]}$ and $m_{k}^{[-]}$ as the number of non-zero elements of the $k$th row of ${\boldsymbol{B}}_{[n-1]}$. We then obtain the bounds
[…]
for every $k$ that does not corresponds to the harmonic component. We recall that we obtain a << instead of the $\leq$ for normal fixed points with $S^{\circ}=\emptyset$. We conclude that we typically need $m_{k}^{[\pm]}\geq 5$ for multistability.
We finally note that the numbers $m_{k}^{[\pm]}$ have an intermediate geometric interpretation: The number $m_{k}^{[+]}$ counts the number of cells of dimension $n+1$ in the boundary of the $k$th cell of dimension $n+2$. The winding number $z_{k}^{[+]}$ counts how often the phases $\boldsymbol{\theta}$$^{[+]}$ wind by $2\pi$ when following all cells in this boundary. Likewise, the number $m_{k}^{[-]}$ counts the number of cells of dimension $n$ that have the $k$th cell of dimension n1n-1 in its boundary.

### v2

@@ Abstract
Higher-order group interactions fundamentally shape the dynamics and stability of oscillator networks. The topological Kuramoto model captures these effects by extending classical synchronization models to include interactions between cells of arbitrary dimension within simplicial and cell complexes. In this article, we present the topological nonlinear Kirchhoff conditions algorithm, a nonlinear generalization of Kirchhoff’s circuit laws, that systematically identifies all phase-locked states in the topological Kuramoto model and reveals how higher-order topology governs multistability. Applying this framework to rings, Platonic solids, and simplexes, we uncover structural cascades of multistability inherited across dimensions, and demonstrate that cell complexes can exhibit richer multistability patterns than simplicial complexes of equal dimension. We find evidence hinting at universal multistability classes determined by the number of boundary cells. These results reveal how higher-order interactions affect synchronization and open new directions for understanding collective dynamics in systems with non-pairwise interactions.
@@ I Introduction
Synchronization phenomena are a hallmark of many natural and engineered systems, from neuronal activity to power grid operation [pikovsky2001synchronization, dorfler2014synchronization]. The Kuramoto model of coupled oscillators has long served as a prototype for understanding how synchronization emerges from local interactions [kuramoto1975self, acebron2005kuramoto]. While early studies focused on globally coupled systems [strogatz2000kuramoto], later extensions to complex networks revealed that network topology crucially shapes both the onset and the nature of synchronization  [rodrigues2016kuramoto]. Such systems often exhibit multistability, where multiple distinct synchronized states coexist [manik2017cycle, delabays2016multistability, jafarpour2022flow, hellmann2020network].
@@
In the classical Kuramoto model, the coupling between the oscillators is pairwise. This assumption is valid for many systems, but in certain contexts where group interactions play an essential role, such as neuronal assemblies [Breakspear2010], the brain connectome [Sizemore2018, Giusti2016, Andjelkovic2020], multimode lasers, [Nixon2013], or protein interaction networks [Estrada2018], it is too restrictive to capture all the relevant couplings in the system. Higher-order connectivity structures that encode such interactions are increasingly recognized as being crucial to many collective dynamics [Millan2025higherorderdynamics, battiston2020networks, battiston2021physics, Bianconi2021book, mulder2018network, Bick2023, Iacopini2019, Santoro2024, Zhang2023HigherOrder]. Recent advances in network science [battiston2020networks, battiston2021physics, Horak2009, Sizemore2018, Petri2013, Petri2014, Giusti2016, Wu2015, Iacopini2019, Andjelkovic2016, Andjelkovic2020, Calmon2022Localdirac, Calmon2022dirac, Calmon2023dirac] have introduced models where the dynamics is not defined only on nodes, but also lives on edges, faces, or higher-dimensional elements of simplicial or cell complexes. Among these, the topological Kuramoto model [millan2020explosive, ghorbanchian2021higher, Nurisso2024UnifiedKuramoto] has become a paradigmatic framework, revealing universal features absent in conventional networks, such as abrupt synchronization transitions  [skardal2020higher, Lucas2020]. Moreover, topologically-induced synchronization obstructions have been found in the study of the simplicial Stuart-Landau model, where global topological synchronization has been studied  [carletti2023global]. While some aspects of multistability [Skardal2019multistability, Deville2021] and phase locking [Nurisso2024UnifiedKuramoto] in similar simplicial models have been analyzed separately, there is still no systematic framework for analyzing how multistable phase-locked states emerge from higher-order interactions.
@@
We address this gap by studying phase locking and multistability in the topological Kuramoto model defined on general cell complexes. Unlike earlier work based on the master stability function formalism  [carletti2023global, Gambuzza2021stability, Parastesh2025nearsynchronization], we develop the topological nonlinear Kirchhoff conditions algorithm, a generalization of Kirchhoff’s current and voltage laws (Fig. 1). Most previous studies have focused on simplicial complexes, yet many real-world structures, such as architectural trusses, social interactions or protein complexes, are better described as cell complexes [hoppe2025don, Roddenberry2022, Giambagli2022, mulder2018network], comprised of larger polyhedral faces. Despite their prevalence, the study of dynamical phenomena on cell complexes is still in its infancy.
Figure 1: From Kirchhoff’s circuit law to the topological nonlinear Kirchhoff conditions. Kirchhoff’s voltage law enforces linear conservation of potential differences around each circuit loop. In the nonlinear generalization, phase differences $\psi_{i}$ across network edges satisfy a nonlinear constraint, $\sum_{i}\arcsin(\psi_{i})=2\pi z$, where integer winding numbers $z$ describe distinct phase-locked states. The topological generalization extends this principle to higher-order interactions on cell complexes, where linear combinations of nonlinear phase differences across higher-dimensional cells satisfy $\sum_{i}C_{i}\arcsin\psi_{i}^{[\pm]}=2\pi z_{i}^{[\pm]}$, with $C_{i}$ determined by the boundary structure of each cell. Here, the dynamics is defined on edges, and interactions occur via nodes in the lower dimension ($\psi_{i}^{[-]}$) and faces in the upper dimension ($\psi_{i}^{[+]}$).
@@
Our algorithm systematically identifies all distinct phase-synchronized states in cell complexes and provides rigorous criteria for when going beyond pairwise interactions pays off in terms of multistability. We demonstrate our approach on three minimal but paradigmatic classes of cell complexes: (i) rings, (ii) Platonic solids, and (iii) simplexes. These basic motifs can be combined to build larger complexes with potentially even richer multistability. Our framework can be applied to any dimension, but we focus on the case most relevant to applications, where the dynamics is defined on edges and interactions occur via nodes and faces [Giambagli2022].
In summary, we show that cell complexes support richer multistability than simplicial complexes of the same dimension. Higher-order connectivity gives rise to structural cascades in which stable states are inherited across dimensions. We also find indications of universality classes due to boundary structure.
@@ II Cell complexes
Cell complexes are generalised network structures that contain cells of different dimensions [hoppe2025don, Sizemore2018, Roddenberry2022]. In addition to nodes $(n=0)$ and edges ($n=1$) found in standard graphs, they can include polygons $(n=2)$, polytopes $(n=3)$ and other higher-dimensional elements. Most studies on synchronization focus on the special case of simplicial complexes  [millan2020explosive, Nurisso2024UnifiedKuramoto], where each cell is an Euclidean simplex, but many real-world systems are better described by cell complexes.
@@
Using this property, one can show that every vector ${\boldsymbol{x}}_{[n]}\in\mathbb{R}^{N_{[n]}}$ can be decomposed according to the Hodge decomposition [barbarossa2020topological] as
@@
Figure 2: Examples of cell complexes for $n=1$. Two representative cell complexes, (a) a ring and (b) a cube, where the dynamics is defined on the edges, and interactions occur via both nodes and faces. Nodes are shown in black, edges in blue and faces in light green.
@@ III Phase locking in the topological Kuramoto model
In this article we consider a generalized topological Kuramoto model [millan2020explosive] allowing for inhomogeneous coupling constants. The basic dynamic variables are the phases $\theta_{i}(t)$ associated to each $n$-dimensional cell. These phases interact via cells of lower and higher dimension. In the case $n=1$ for example, phases are defined for each edge, and interact via vertices and faces. Summarizing the dynamical variables in the vector ${\boldsymbol{\theta}}=\left(\theta_{1},\ldots\theta_{N_{[n]}}\right)^{\top}$, the equations of motion read
@@
We can decouple the dynamics by multiplying the equations of motion by ${\boldsymbol{B}}_{[n]}$ and ${\boldsymbol{B}}_{[n+1]}^{\top}$ from the left [millan2020explosive]. Defining the variables
@@
In general, a stationary state is linearly stable if the real part of all eigenvalues is negative [strogatz2018nonlinear]. If the stationary state satisfies the conditions
@@
we find that all eigenvalues are real and satisfy the following conditions: (i) If an eigenvector $\boldsymbol{v}$} is purely harmonic, then the associated eigenvalue is given by $\lambda=0$. (ii) If an eigenvector $\boldsymbol{v}$} is not purely harmonic, then the associated eigenvalue satisfies $\lambda<0$ (see appendix for a proof). We thus conclude that a stationary state that satisfies Eq. (11) is linearly stable with respect to perturbations that are orthogonal to the harmonic subspace and neutrally stable with respect to harmonic perturbations. Since harmonic perturbations do not affect the projected quantities $\boldsymbol{\theta}$$^{[+]}$ and $\boldsymbol{\theta}$$^{[-]}$, phase locking is linearly stable.
In the following we will mainly concentrate on phase locked states that satisfy the conditions in Eq. (11) and refer to them as normal fixed points. Stationary states that violate Eq. (11) are typically, but not always, linearly unstable. For the ordinary Kuramoto model $n=0$, it has been shown that stable phase locked states that violate these conditions exist at the border of the stability region [manik2014supply]. We will return to this question later in the context of polygonal face complexes, where we find such stable phase-locked states that aren’t normal fixed points (Fig.3).
IV The topological nonlinear Kirchhoff conditions
@@
In the case $n=1$, the original dynamical variables $\boldsymbol{\theta}$} are defined on the edges of the cell complex, while the variables $\boldsymbol{\psi}$$^{[+]}$ are defined on the faces, and the variables $\boldsymbol{\psi}$$^{[-]}$ on the nodes. In the case $n=0$, the original dynamical variables $\boldsymbol{\theta}$} are defined on the nodes of the graph, while the auxiliary variables $\boldsymbol{\psi}$$^{[+]}$ are defined on its edges. In the theory of supply networks, the entries of ${\boldsymbol{K}}_{[1]}{\boldsymbol{\psi}}^{[+]}$ are interpreted as flows [hartmann2024synchronized, Bacic2025phasecohesive, Parameswaran2025symmetrybreaking].
@@
The complexity of finding a stationary state is then shifted to the question of whether a solution candidate (16) exists that can be expressed as (12). To approach this topic, we first discuss how to recover the original variables $\boldsymbol{\theta}$$^{[\pm]}$ from the $\boldsymbol{\psi}$$^{[\pm]}$. If the equations Eq. (12) have solutions, they can have the form $\arcsin(\cdot)$ or $\pi-\arcsin(\cdot)$.
@@
$f^{\bullet}(\psi_{i}^{[\pm]})=\arcsin(\psi_{i}^{[\pm]})$ (19)
up to integer multiples of $2\pi$ and collect the respective cells in the sets $S_{[n+1]}^{\bullet}$ and $S_{[n-1]}^{\bullet}$. Likewise, if the cosine is not positive, we define the inverse
$f^{\circ}(\psi_{i}^{[\pm]})=\pi-\arcsin(\psi_{i}^{[\pm]})$ (20)
up to integer multiples of $2\pi$ and collect the respective cells in the sets $S_{[n+1]}^{\circ}$ and $S_{[n-1]}^{\circ}$. Finally, we summarize the resulting components in the two vector-valued functions ${\boldsymbol{f}}_{[n+1]}({\boldsymbol{\psi}}^{[+]})$ and ${\boldsymbol{f}}_{[n-1]}({\boldsymbol{\psi}}^{[-]})$.
@@
where $\boldsymbol{z}$$^{[\pm]}$ is a vector of integers. In the case of ordinary graphs, the vectors $\boldsymbol{z}$$^{[\pm]}$ are referred to as winding vectors and its components as winding numbers [delabays2016multistability, manik2017cycle, jafarpour2022flow]. We will adopt this notation in the following and discuss the interpretation below.
@@
We briefly discuss three important aspects of this algorithmic approach to phase locking. First, phase locked states with $S^{\circ}=\emptyset$ are always linearly stable and as such are especially relevant for applications. States with $S^{\circ}\neq\emptyset$ are typically, but not always unstable. In fact, we will demonstrate the existence of stable phase locked states with $S^{\circ}\neq\emptyset$ close to bifurcations (cf. Fig. 4).
Second, the algorithm can be viewed as a topological, nonlinear generalization of Kirchhoff’s circuit laws (Fig. 1). This is most easily see for the case of ordinary graphs $n=0$. Here, the entries of ${\boldsymbol{K}}_{[1]}{\boldsymbol{\psi}}^{[+]}$ are interpreted as flows or currents defined on edges [hartmann2024synchronized, Bacic2025phasecohesive, Parameswaran2025symmetrybreaking]. The set of solution candidates obtained by Eq. (16) correspond to all states that satisfy Kirchhoff’s current law at each node. These states can be decomposed into a special solution plus a set of loop flows. If we linearize the sine and arcsine functions and set ${\boldsymbol{z}}^{[\pm]}={\boldsymbol{0}}$, then conditions (22) correspond to Kirchhoff’s voltage law. Together, Kirchhoff’s voltage and current laws then uniquely determine the currents.
Algorithm 1 introduces two generalizations: (i) The nonlinearity of the sine introduces multistability. Since $\sin(\alpha)=\sin(\alpha+2\pi z)$ for all $z\in\mathbb{Z}$, we can obtain different values for phase variables when inverting the sine function. Hence, we can obtain different solutions parametrized by the winding numbers $z$ [delabays2016multistability, manik2017cycle, jafarpour2022flow]. (ii) The algorithm generalizes the nonlinear Kirchhoff conditions to cell complexes of arbitrary dimensions.
Third, the possible values of the entries of the vectors $\boldsymbol{z}$$^{[\pm]}$ and thus the number of steady states is bounded depending on the topology of the cell complex. This bound is essentially determined by the number of of cells of higher or lower dimension that are adjacent to a given cell $k$. More precisely, let $m_{k}^{[+]}$ denote the number of cells of dimension $n+1$ in the boundary of the $k$th cell of dimension $n+2$ and $m_{k}^{[-]}$ the number of cells of dimension $n$ that have the $k$th cell of dimension n1n-1 in its boundary. We then obtain the bounds
@@
for every $k$ that does not corresponds to the harmonic component (see appendix for a proof). In the case of normal fixed points with $S^{\circ}=\emptyset$ we obtain a << instead of the $\leq$. We conclude that we typically need $m_{k}^{[\pm]}\geq 5$ for multistability.